# A quantitative analysis of semantic information in deep representations of text and images

### v3

@@ 5 Conclusion
Through a novel application of the information imbalance measure II , we showed that DeepSeek-V3, the largest publicly available LLM, developed an internal processing phase in which different inputs that share the same semantics, such as translations and image-caption pairs, are reflected in representations that are extremely similar. We further ascertained that, on these “semantic” layers, long token spans meaningfully contribute to the representation, and long-distance correlations in encoded information cue high-quality representations. Several of the patterns we observed in DeepSeek-V3 also emerge in the medium-sized Llama3.1-8b model, but in a less distinct manner, suggesting that deep semantic processing is a hallmark of better models. We also analyzed visual processing models, finding again the presence of layers that capture deeper semantic similarity, and whose position depends on the objective the models are trained upon. DeepSeek-V3 textual representations and vision transformer visual representations of the same concepts are most strongly aligned in the respective semantic layers.
Our work supports the hypothesis that, as deep models improve, they converge towards shared representations of the world platonic_hypothesis . We took a first step away from simply verifying that such representations exist, towards characterizing their nature, by precisely quantifying the degree of shared information, localizing where they occur in different networks and determining how they are synthesized from composite inputs. Future research, besides extending the empirical coverage of our investigation, should further increase the resolution of the measurements, allowing us to go from observing the holistic behavior of a model with respect to a data set to a full characterization of how semantic representations are constructed given each specific input instance.
@@ Appendix A Compute resources
We run DeepSeek-V3 on a cluster of 16 H100 GPUs (80GB each), using the SGLang framework [41]. Input sequences consist of 40–80 tokens and are processed in batches of 30. For LLama3.1-8B, a single GPU from a DGX H100 system is sufficient. For our experiments on text and image-caption pairs, 1TB of RAM is enough, moving to the GPU a couple of layers at a time. Image-Image experiments are done using up to five NVIDIA A30 GPUs and 500 GB of RAM.
@@ Appendix B Assets
Appendix C Information Imbalance: implementation details and Gaussian comparisons
Implementation details
The representations of deep transformers are (very) high dimensional: the number of tokens ranges from a few tens for short sentences, to thousands for long sentences or images, and typical (per token) embedding dimensions are of the order $\mathcal{O}(10^{3})$. These two contributions rapidly generate representations where the total dimensionality can be of order $\mathcal{O}(10^{6})$. Estimating the information imbalance in such a large space can be very costly, especially with sample sizes of order $10^{3}$, like those used in this work. To speed up computations, we follow the approach introduced in Ref. [42]. Instead of directly computing all-to-all distances between the $N$ samples, we randomly split the data, selecting two subsets of points, that we call origins and destinations. We call $N_{o}$ and $N_{d}$ the number of samples in both subsets, respectively. Computing the information imbalance without splitting is costly for two reasons. First, it is necessary to compute the pairwise distances between all $N$ data points, instead of using only the $N_{o}$ origins. Second, the ranking operations sort N1N-1 elements instead of $N_{d}-1$ elements, which strongly affects compute time, as sorting operations grow more than linearly with the number of elements. In this work, we found a good number of origin points to be $\approx 200-500$. Indeed, with these numbers the statistical uncertainty on the II is small enough to assess subtle differences between the languages, as shown in the figures in the main text (notice that the error bars are smaller than the marker size in all the figures except in Fig. 4, where the error bars are explicitly shown).
Interpreting II
To provide an interpretation of the value of the information imbalance observed in our analysis of the representations of images and text, we computed the II in an idealized scenario, in which the dimensions of the feature spaces resemble the one in neural networks: we generate set $Y$ with 2000 samples from a Gaussian distribution with zero-mean, identity variance matrix, and $p=10^{5}$ components. We fix $X$ to be the first $k$ components of $Y$, with $k=1,2,\dots,10^{5}$. We compute the information imbalance $\Delta\left(X\to Y\right)$ to quantify how adding features changes the II. We repeat the analysis by binarizing the features, namely following the same procedure used for the activations in the neural networks. In the real case, we use the $L^{2}$ distance; in the binary case we employ the Hamming distance. We report the results in Fig.5: we average the results over 10 replicas. We see that the information imbalance between $X$ and $Y$ decreases as the number of shared features increases in a highly nonlinear manner, and that the results with and without binarization are within statistical error. In Sec. H we further show that binarization also preserves the Information Imbalance calculated with activations generated by DeepSeek-V3.
Figure 5: Information imbalance for growing shared components in simulated data. We compute the information imbalance $\Delta(X\to Y)$, where $Y$ is the target space and $X$ is the origin space: it consists of the first $k$ features of $Y$, with $k=1,\dots,10^{5}$. We report both the real (Euclidean distance) and the binary cases (Hamming distance). The results are averaged across ten repetitions; the shades are the 1-sigma bootstrap errors.
Appendix D Translations dataset
@@
We filter pairs of translated sentences from Helsinki-NLP/opus_books [33] that have a number of tokens between $40$ and $80$, avoiding trivial sentences and incorrect translations, while keeping roughly 5000 thousands sentences of comparable length. We avoid truncating the translated sentences, preserving their full meaning. We excluded the last two tokens of each sentence from the analysis, as they consistently correspond to punctuation marks (e.g., a period or a period followed by a quotation mark), which introduce trivial similarities in the representations.
Appendix E Average token comparison
@@
Fig. 6 shows the Information Imbalance between DeepSeek-V3 representations of English and Spanish translations, using 20 tokens of each sentence, with and without taking the average over the token axis. Note that for the case of concatenated tokens, the curve is the same as in Fig. 1 of the main article, for direct comparison. When taking the average, we use the $L_{2}$ distance and we clip and normalize the inputs, as Ref. [3]. When using all 20 tokens we binarize activations and we use the Hamming distance, as in the main text. While the results are qualitatively consistent, they differ significantly in quantitative terms. Across all layers—most notably in the first and last—averaging leads to a marked increase in mutual predictability (lower Information Imbalance), due to the elimination of positional information, which is most relevant in the initial and last layers, i.e., before and after the semantic region.
@@
Appendix F Misalignment of translations erases semantic similarity
As a consistency check, Fig 7 shows the Information Imbalance for DeepSeek-V3’s and Llama3.1-8b representations using misaligned translations, namely performing a batch-shuffle in one of the datasets. Since the semantic correspondence between sentences is destroyed, the representations are not informative about each other, and thus the II is close to one, for all layers. The same occurs with misaligned image pairs, II being around one when pairs do not share the same class, for all layers.
@@ Appendix F Misalignment of translations erases semantic similarity
Appendix G Comparison between neighborhood overlap and Information Imbalance
@@
To quantify representation similarity, Ref. [3] used the neighborhood overlap, namely the average fraction of shared $k$ nearest neighbors. In particular, they found this metric to be more suitable than linear metrics, like Central Kernel Alignment (CKA), providing a stronger signal. In this section we compare our results from Fig. 1.a) of the main article with the neighborhood overlap computed on the same representations. To measure the neighborhood overlap we follow the pipeline of Ref. [3], namely we clip the activations using quantiles of order $0.05\%$ and $95\%$, we normalize each vector, and we measure the distances between them with the $L_{2}$ metric. Fig 8 shows that the neighborhood overlap between representations increases and reaches a plateau concurrently with the local minimum plateau of the Information Imbalance, thus both methods give qualitative the same results. We note that the neighborhood overlap, being defined between 0 and 1, here only reaches a value around 0.2. Similar small values were also observed and even highlighted in Ref. [3], although the authors do not provide arguments to understand the reasons behind this property. We note that if the $k$-th neighbor of representation $A$ is instead the $k+1$ neighbor of representation $B$, automatically that point counts as outside the $k$-neighborhood, even if it is extremely close, rendering the alignment value very low. This effect doesn’t take place in the Information Imbalance, which, instead of evaluating if $k$ neighbors are shared, measures what is the rank in space $B$ of $k$ neighbors in space $A$, and vice-versa.
@@
Figure 8: Information Imbalance between English (en) and Spanish (es) representations generated by DeepSeek-V3 from Fig. 1 of the main article, compared with the neighborhood alignment metric from Ref. [3]. The (hardly visible) shaded area corresponds to one standard deviation, computed with a Jackknife procedure subsampling 2000 samples out of 5000.
@@
Appendix H Binarizing high dimensional representations has a marginal effect on the Information Imbalance
@@
Fig. 9 includes again, as reference, the Information Imbalance between DeepSeek-V3 representations of English and Spanish translations, using the last 20 tokens of each sentence, binarizing the activations and using the Hamming distance, (results of Fig. 1 from main text). To investigate the effects of binarizing the representations, we show the Information Imbalance using the same data where, instead of taking the sign of the activations to generate binary variables, we clip them using quantiles of order $0.05\%$ and $95\%$, we normalize vectors to unit norm, and we use the $L_{2}$ distance, similar to Ref. [3]. Remarkably, binarizing has marginal effects on the Information Imbalance computed with full precision (for DeepSeek-V3’s representations, BF16), which is possible in this setup given that we are working with 20 tokens.
@@
Figure 9: In circles, Information Imbalance between English (en) and Spanish (es) representations generated by DeepSeek-V3, binarizing the activations and using the Hamming distance. In crosses, the II clipping and normalizing activations, and using the $L_{2}$ metric. The standard deviation, computed with a Jackknife procedure subsampling 2000 samples out of 5000, is smaller than marker size.
Appendix I Binary Intrinsic Dimension of DeepSeek-V3 representations
@@
The Binary Intrinsic Dimension (BID) is defined in Ref. [28] through the probability distribution of Hamming distances between samples of binary high dimensional vectors. If we consider $N$-dimensional vectors $\bm{\sigma}$ with components $\sigma_{i}$, i=1,,Ni=1,...,N uniformly distributed in $\{0,1\}$, then the probability of observing Hamming distance $r$ between any two samples is exactly $P_{0}(r)=\frac{1}{2^{N}}\binom{N}{r}$. The BID is defined as the coefficient $d_{0}$ of the following Ansatz:
@@
where $d_{1}$ is a second variational parameter. We perform a second-order optimization of the Kullback-Leibler divergence between the model equation 2 and the empirical distribution of distances, computing numerical derivatives with JAX’s autodiff. For more details on the BID computation, see [28]. We used hyperparameters $\alpha_{min}=0.15$, and $\alpha_{max}=0.4$. Fig. 10 shows the Binary Intrinsic Dimension (BID) of DeepSeek-V3’s representations processing English sentences from opus_books with length between 40 and 50 tokens. We take the last 10 and the last 20 tokens of the binarized representation, and we find two dimensionality peaks that roughly coincide with the beginning (around 0.4) and the end (around 0.9) of the semantic region found in Fig. 1. These results align with recent evidence of abstract spaces presenting dimensionality peaks in deep representations [17, 32]. The missing points at the beginning of the blue curve correspond to failed optimizations due to the strong multimodality of the empirical distribution of distances.
@@
Appendix J Image-Captions Pairs
In Fig. 11, we report additional results for the image-caption experiments: we consider different windows of tokens, using the last 200 tokens for images - as we do in the main text - vs a smaller windows of 50 tokens and the average. In general, results are qualitatively similar across setups. We observe that the choice of the number of tokens comes with a trade-off: more tokens convey more information; at the same time, having more tokens implies working with bigger objects.

### v4

@@ 5 Conclusion
Through a novel application of the Information Imbalance measure [II], we showed that DeepSeek-V3, the largest publicly available LLM, developed an internal processing phase in which different inputs that share the same semantics, such as translations and image-caption pairs, are reflected in representations that are extremely similar. We further ascertained that, on these “semantic” layers, long token spans meaningfully contribute to the representation. Several of the patterns we observed in DeepSeek-V3 also emerge in the medium-sized Llama3.1-8b model, but in a weaker manner, suggesting that deep semantic processing is a hallmark of better models. We also analyzed visual processing models, finding again the presence of layers that capture deeper semantic similarity, and whose position depends on the objective the models are trained upon. DeepSeek-V3 textual representations and vision transformer visual representations of the same concepts are most strongly aligned in the respective semantic layers.
Our work supports the hypothesis that, as deep models improve, they converge towards shared representations of the world [platonic_hypothesis]. We took a first step away from simply verifying that such representations exist, towards characterizing their nature, by precisely quantifying the degree of shared information, localizing where they occur in different networks and determining how they are synthesized from composite inputs. Future research, besides extending the empirical coverage of our investigation, should further increase the resolution of the measurements, allowing us to go from observing the holistic behavior of a model with respect to a data set to a full characterization of how semantic representations are constructed given each specific input instance.
@@ Appendix A Compute resources
We run DeepSeek-V3 on a cluster of 16 H100 GPUs (80GB each), using the SGLang framework [zheng2024sglang]. For LLama3.1-8B, a single GPU from a DGX H100 system is sufficient. For our experiments on text and image-caption pairs, 1TB of RAM is enough, moving to the GPU a couple of layers at a time. Image-Image experiments are done using up to five NVIDIA A30 GPUs and 500 GB of RAM.
@@ Appendix B Assets
DeepSeek-V3: https://huggingface.co/deepseek-ai/DeepSeek-V3 ; Model licence.
Llama3.2-1b: https://huggingface.co/meta-llama/Llama-3.2-1B ; license: llama3.2
Llama3.2-3b: https://huggingface.co/meta-llama/Llama-3.2-3B ; license: llama 3.2
@@
BERT: https://huggingface.co/google-bert/bert-base-multilingual-cased ; licence: apache-2.0
@@
Appendix C Translations dataset
@@
We filter pairs of translated sentences from Helsinki-NLP/opus_books [opus_books] that have a number of tokens between $40$ and $80$, avoiding trivial sentences and incorrect translations, while keeping roughly 5000 thousands sentences of comparable length. We avoid truncating the translated sentences, preserving their full meaning. We excluded the last two tokens of each sentence from the analysis, as they consistently correspond to punctuation marks (e.g., a period or a period followed by a quotation mark), which introduce trivial similarities in the representations.
Appendix D Average token comparison
@@
Fig. 5 shows the Information Imbalance between DeepSeek-V3 representations of English and Spanish translations, using 20 tokens of each sentence, with and without taking the average over the token axis. Note that for the case of concatenated tokens, the curve is the same as in Fig. 2 of the main article, for direct comparison. When taking the average, we use the $L_{2}$ distance and we clip and normalize the inputs, as platonic_hypothesis. When using all 20 tokens we binarize activations and we use the Hamming distance, as in the main text. While the results are qualitatively consistent, they differ significantly in quantitative terms. Across all layers—most notably in the first and last—averaging leads to a marked increase in mutual predictability (lower Information Imbalance), due to the elimination of positional information, which is most relevant in the initial and last layers, i.e., before and after the semantic region.
@@
Appendix E Misalignment of translations erases semantic similarity
As a consistency check, Fig 6 shows the Information Imbalance for DeepSeek-V3’s and Llama3.1-8b representations using misaligned translations, namely performing a batch-shuffle in one of the datasets. Since the semantic correspondence between sentences is destroyed, the representations are not informative about each other, and thus the Information Imbalance is close to one, for all layers. The same occurs with misaligned image pairs, Information Imbalance being around one when pairs do not share the same class, for all layers.
@@ Appendix E Misalignment of translations erases semantic similarity
Appendix F Comparison between neighborhood overlap and Information Imbalance
@@
To quantify representation similarity, platonic_hypothesis used the neighborhood overlap, namely the average fraction of shared $k$ nearest neighbors. In particular, they found this metric to be more suitable than linear metrics, like Central Kernel Alignment (CKA), providing a stronger signal. In this section we compare our results from Fig. 1.a) of the main article with the neighborhood overlap computed on the same representations. To measure the neighborhood overlap we follow the pipeline of platonic_hypothesis, namely we clip the activations using quantiles of order $0.05\%$ and $95\%$, we normalize each vector, and we measure the distances between them with the $L_{2}$ metric. Fig 7 shows that the neighborhood overlap between representations increases and reaches a plateau concurrently with the local minimum plateau of the Information Imbalance, thus both methods give qualitative the same results. We note that the neighborhood overlap, being defined between 0 and 1, here only reaches a value around 0.2. Similar small values were also observed and even highlighted in platonic_hypothesis, although the authors do not provide arguments to understand the reasons behind this property. We note that if the $k$-th neighbor of representation $A$ is instead the $k+1$ neighbor of representation $B$, automatically that point counts as outside the $k$-neighborhood, even if it is extremely close, rendering the alignment value very low. This effect doesn’t take place in the Information Imbalance, which, instead of evaluating if $k$ neighbors are shared, measures what is the rank in space $B$ of $k$ neighbors in space $A$, and vice-versa.
@@
Figure 7: Information Imbalance between English (en) and Spanish (es) representations generated by DeepSeek-V3 from Fig. 2 of the main article, compared with the neighborhood alignment metric from platonic_hypothesis. The (hardly visible) shaded area corresponds to one standard deviation, computed with a Jackknife procedure subsampling 2000 samples out of 5000.
@@
Appendix G Binarizing high dimensional representations has a marginal effect on the Information Imbalance
@@
Fig. 8 includes again, as reference, the Information Imbalance between DeepSeek-V3 representations of English and Spanish translations, using the last 20 tokens of each sentence, binarizing the activations and using the Hamming distance, (results of Fig. 2 from main text). To investigate the effects of binarizing the representations, we show the Information Imbalance using the same data where, instead of taking the sign of the activations to generate binary variables, we clip them using quantiles of order $0.05\%$ and $95\%$, we normalize vectors to unit norm, and we use the $L_{2}$ distance, similar to platonic_hypothesis. Remarkably, binarizing has marginal effects on the Information Imbalance computed with full precision (for DeepSeek-V3’s representations, BF16), which is possible in this setup given that we are working with 20 tokens.
@@
Figure 8: In circles, Information Imbalance between English (en) and Spanish (es) representations generated by DeepSeek-V3, binarizing the activations and using the Hamming distance. In crosses, the Information Imbalance clipping and normalizing activations, and using the $L_{2}$ metric. The standard deviation, computed with a Jackknife procedure subsampling 2000 samples out of 5000, is smaller than marker size.
Appendix H Binary Intrinsic Dimension of DeepSeek-V3 representations
@@
The Binary Intrinsic Dimension (BID) is defined in acevedo2025 through the probability distribution of Hamming distances between samples of binary high dimensional vectors. If we consider $N$-dimensional vectors $\bm{\sigma}$ with components $\sigma_{i}$, i=1,,Ni=1,...,N uniformly distributed in $\{0,1\}$, then the probability of observing Hamming distance $r$ between any two samples is exactly $P_{0}(r)=\frac{1}{2^{N}}\binom{N}{r}$. The BID is defined as the coefficient $d_{0}$ of the following Ansatz:
@@
where $d_{1}$ is a second variational parameter. We perform a second-order optimization of the Kullback-Leibler divergence between the model equation 1 and the empirical distribution of distances, computing numerical derivatives with JAX’s autodiff. For more details on the BID computation, see acevedo2025. We used hyperparameters $\alpha_{min}=0.15$, and $\alpha_{max}=0.4$. Fig. 9 shows the Binary Intrinsic Dimension (BID) of DeepSeek-V3’s representations processing English sentences from opus books with length between 40 and 80 tokens. We take the last 10 and the last 20 tokens of the binarized representation, and we find two dimensionality peaks that roughly coincide with the beginning (around 0.4) and the end (around 0.9) of the semantic region found in Fig. 2. These results align with recent evidence of abstract spaces presenting dimensionality peaks in deep representations [cheng2025emergence, antonello2024evidence]. The missing points at the beginning of the blue curve correspond to failed optimizations due to the strong multimodality of the empirical distribution of distances.
@@
Appendix I Token-Level Information Imbalance and Causal Asymmetry Across Languages
Figure 10: Panel a) In solid markers, the “backward” Information Imbalance, i.e., from the last token to a previous token at token-distance $\tau$, as a function of $\tau$, for DeepSeek-V3 representations at layer 43 and sentences in several languages. In dashed lines, the reciprocal “forward” Information Imbalance. Panel b) Asymmetry $\mathcal{A}$ between forward and backward Information Imbalance, averaged over $\tau$, as a function of DeepSeek’s relative depth. The Jackknife error bars (five repetitions subsampling 1,000 of 2,000 sentences) are too small to show; the marker sizes act as an upper bound on these errors.
Fig. 10a) focuses on a semantic layer from DeepSeek-V3, and it shows token-token Information Imbalance as a function of token distance for several languages. We present the Information Imbalance in both directions: The "backward" Information Imbalance, from the last token, $T$, to a previous token at distance $\tau$, $\Delta(T\rightarrow T-\tau)$, and the reciprocal or “forward” Information Imbalance, $\Delta(T-\tau\rightarrow T)$. We observe that English deep representations are much more correlated than the representations in the other languages. We conjecture that this is a sign of better quality, consistent with Fig. 2b), where the presence of English in a pair results in the best predictability (lowest Information Imbalance scores). At the opposite extreme, note that single-language German and Hungarian representations are the least correlated in Fig. 10a), and the German-Hungarian translations have indeed the worst mutual predictability score (highest Information Imbalance) in Fig. 2. Furthermore, Spanish and Italian have the second and third most correlated inner representations in Fig. 10a), coherently with the profiles of the English-Spanish and English-Italian pairs in Fig. 2b) (lowest and second lowest Information Imbalance values, respectively).
Finally, Fig. 10a) shows that, for English, as the distance between tokens $\tau$ increases, the relative information content between past and future tokens gets increasingly asymmetric. In particular, earlier tokens are more informative of the last one than the other way around. This could be expected, since causal models like DeepSeek-V3 are trained to predict future tokens, and text itself has a natural causal asymmetry. Nonetheless, Fig. 10a) shows that this past-to-future asymmetry is heterogeneous across languages. To further study the effect, we first define the asymmetry as $\mathcal{A}=\Delta(T\rightarrow T-\tau)-\Delta(T-\tau\rightarrow T)$. When $\mathcal{A}$ is positive, the earlier token predicts the last one more than the latter predicts the earlier token, whereas a negative $\mathcal{A}$ has the opposite interpretation. Fig. 10b) shows the information asymmetry $\mathcal{A}$ of DeepSeek-V3 as a function of its layers for the all languages we studied. For initial layers (i.e., relative depth less than roughly $0.4$, see Fig. 2), we observe that all representations have a comparable amount of causal asymmetry. Instead, for semantic layers there are strong heterogeneities between languages. English remains causally asymmetric until the last few layers of the network, while the other languages present earlier negative values of $\mathcal{A}$. Intriguingly, Spanish and Italian, the two languages with the lowest Information Imbalance values from English, have a later dip in $\mathcal{A}$ compared to German and Hungarian. These results call for further experiments to study possible relationships between these observed information asymmetries, model performance, and linguistic structure. We leave them as material for future research. In Sec. J we study this information asymmetry for a non autoregressive model.
Appendix J BERT’s forward-backward asymmetry profile
In this section we study, the asymmetry in II between two tokens at distance $\tau$, $\mathcal{A}=\Delta(T\rightarrow T-\tau)-\Delta(T-\tau\rightarrow T)$, where $T$ is the last token, to see if the observed asymmetries are similar to those of DeepSeek-V3 (Fig. 10b). Fig. 11 shows the asymmetry $\mathcal{A}$ for similar English sentences from opus books as those used in Fig. 10b), processed by BERT. Contrary to what was observed in Fig. 10b), BERT’s representations present a positive asymmetry close to the initial and final layers, and negative asymmetry in central layers. This strong difference of asymmetry profiles opens the door for further studies on the asymmetry of correlations in several different models and architectures, and its potential relations to downstream tasks.
Figure 11: Asymmetry $\mathcal{A}$ between forward and backward Information Imbalance, averaged over token all distances as a function of BERT’s relative depth. The shaded area corresponds to one standard deviation computed subsampling half out of 7500 samples, five times.
Appendix K Image-Captions Pairs
In Fig. 12, we report additional results for the image-caption experiments: we consider different windows of tokens, using the last 200 tokens for images –as we do in the main text– vs a smaller windows of 50 tokens and the average. In general, results are qualitatively similar across setups. We observe that the choice of the number of tokens comes with a trade-off: more tokens convey more information; at the same time, having more tokens implies working with bigger objects.